# ExpliCa: Evaluating Explicit Causal Reasoning in Large Language Models

### v3

@@ 4.1 Models
We selected $7$ generative LLMs. Specifically, we selected two open-weights models (Mistral-7B-Instruct-v0.3 and falcon-7b-instruct), three partially open models (Meta-Llama-3.1-8B-Instruct, gemma-2-9b-it, and Qwen2.5-7B-Instruct), and two commercial models, gpt4o and gpt4o-mini. Perplexity evaluation was not performed on commercial models as it is not permitted through the API. We used Qwen2.5 instruct models of different sizes (from $0.5$B to $32$B parameters) for analyzing the impact of model scale. [^13] More details on models in App. G.
@@ 5 Results and Discussion
We present the results of our experiments by first providing an overview of the core outcomes from the prompting- and perplexity-based evaluations (Sec. 5.1). Then, we provide an in-depth analysis of the top-performing models, and a comparative analysis of Qwen models across different sizes (in Sec. 5.2). The fine-grained results, including performances of each model, task, relation type, and setting are reported in Appendix H.
5.1 Core Analysis Results
The core analysis of the models on the ExpliCa benchmark primarily involves comparing their linguistic performance through prompting and evaluating their competence in causal reasoning using perplexity. We also provide a comprehensive overview of the results, organized by both task and model, revealing substantial variability in performance across the prompting tasks.
Modeling PCD with Prompting.
Figure 3: Models’ results with APS scores (top, bottom), Greedy Search (top), and Outlines (bottom).
@@ 5.1 Core Analysis Results
Acc. $0.49$±$\pm 0.2$ $0.43$±$\pm 0.2$ $0.66$±$\pm 0.0$ $0.40$±$\pm 0.2$ $0.32$±$\pm 0.1$ $0.62$±$\pm 0.2$
Cloze $0.38$±$\pm 0.1$ $0.32$±$\pm 0.1$ $0.54$±$\pm 0.0$ $0.40$±$\pm 0.1$ $0.35$±$\pm 0.1$ $0.53$±$\pm 0.0$
M.C. $\pm 0.2$ $0.18$±$\pm 0.2$ $0.27$±$\pm 0.3$ $0.38$±$\pm 0.1$ $0.34$±$\pm 0.1$ $0.46$±$\pm 0.2$
Avg $0.36$±$\pm 0.2$ $0.31$±$\pm 0.2$ $0.49$±$\pm 0.2$ 0.39±$\pm 0.2$ $0.34$±$\pm 0.1$ $0.54$±$\pm 0.2$
Fig. 3 shows results for each model in each task in both greedy search and Outlines settings. APS scores are reported in both charts for comparison. GPT-4o is the best-performing model in most tasks. Higher results are obtained mostly by models in the acceptability rating task (zero or few-shot), with a few exceptions, like Mistral. Moreover, the Outlines framework proved to be beneficial for all models on the multiple-choice task. Conversely, for the cloze test and acceptability rating task, performances either remain roughly the same or decrease substantially for both open models and GPTs. Notably, GPT4o-mini exhibits a much wider performance range across both few-shot and zero-shot settings. Moreover, each model performs best under different settings, in particular GPT4o, Gemma, and Qwen. GPT4o seems to benefit from few-shot learning in all tasks and decoding strategy, more than the other models.
Modelling PCD with Perplexity.
The APS results of the open models are shown in Fig. 3. The scores obtained by relying solely on perplexity are significantly higher ($0.63$ overall average) compared to those achieved in the prompting tasks. [^14] Perplexity values are in Tab. 9, App. H. Although Falcon performs poorly in the latter, it is the model with the best APS ($0.66$). By leveraging perplexity, the lowest APS too, reached by Qwen with $0.59$, surpasses the other open models in prompting tasks and GPT4o-mini as well. Although this result does not match GPT4o performances, it seems to confirm that models’ competence about causal relations encoded in their probabilistic predictions is more accurate than their prompting performances Hu and Levy (2023).
@@
Overview.
Tab. 2 shows results aggregated by prompting task. On average, PCD performed best when responses were framed using Outlines in a few-shot setting ($0.40$ overall, $0.61$ for GPT models). However, the highest task-specific accuracy was achieved in the acceptability rating task in a zero-shot setting with greedy search ($0.49$). Standard deviation analysis (sd) indicates that task selection significantly affects smaller models’ performance, especially with greedy search (sd $0.2$). Strikingly, the top average accuracy of $0.61$ still reveals a strong gap with human data.
Core results by model are summarized in Fig. 4, [^15] Detailed results are shown in Tab. 6, App. H. which shows the average accuracy across all three prompting tasks for a specific model in a specific setting (e.g., Outlines in zero-shot). Violet bars represent the APS (a single numerical value) for each open model.
We observe a substantial variability of performances even among models of the same scale in prompting tasks, especially for open models with greedy search. Besides the high degree of variation, two trends can be observed: i.) GPT4o outperforms all models, but its mini variant does not follow this trend and is still far from human performance; ii.) Prompting performance is systematically lower than competence-level APSs in open models, which is stable across models.
5.2 In-depth Analysis Results
@@
First, we focus on the relation type between the events expressed in each dataset item, and their order. We report results on the acceptability rating prompting task, with zero-shot and greedy search (i.e., where the models performed best overall). [^16] All the models’ results, in all settings, are in App. H. We specifically focus on GPT4o and two open models, namely Gemma and Falcon, with the latter achieving the best APS. In Fig. 5, we see that models perform best on the causal relations in iconic order, with the exception of Falcon. [^17] Values are shown in Tab. 10, App. H. Worse performances are obtained for causal anti-iconic and temporal iconic, with the latter often being mistaken for causal iconic.
@@
An interesting pattern emerges by comparing APSs and Accuracy on acceptability ratings. APSs for a model are always higher than the respective acceptability ratings’ accuracy, except for temporal anti-iconic items. [^18] As for most of the models (see Tab. 9, App. H). However, by observing the models’ behavior, [^19] See App. I, Fig. 13 for the confusion matrices. we saw that both Gemma and Falcon tend to favor the temporal anti-iconic connective despite showing lower perplexity for the connective expressing a causal iconic relation (i.e., so). GPT4o scores do not seem to suffer the same biases toward specific cases. Nevertheless, intra-cases differences are up to 26%.
@@
We compared human ratings with model-generated ones in terms of their distribution and correlation. We show a comparison of models acceptability ratings with humans’ for the causal-temporal and unrelated subsets in Fig. 6.202020Answers without a rating are not shown in the distribution. We consider ratings of GPT4o and Gemma, [^21] Results for the zero-shot prompt with greedy search. These are the models with the highest accuracy for this task. and the (normalized) perplexity scores for Falcon, as it obtained the best APSs. A noticeable difference exists between human ratings of causal and temporal items vs. unrelated and model scores. Human ratings are characterized by a clear distinction between the relation types, with causal-temporal judgments mostly falling in the 4–8 range and unrelated judgments never exceeding 6 (on a 1–10 scale). There is only a small overlap between the two distributions in the human ratings. However, this is not true for models: GPT4o judgments for causal-temporal are more dispersed along the whole scale, whereas unrelated pairs present many outliers in the top end of the distribution; in Gemma’s ratings there is a larger overlap between causal-temporal and unrelated items; Falcon’s perplexity scores exhibit a similar overlap between the two distribution, but are also more compressed in the lower end of the distribution.
Fig. 7 shows Spearman correlation computed among human ratings and models’ acceptability ratings collected in zero-shot scenario. [^22] Individual results in Tab. 14, App. J. We observed that even though GPT4o, Gemma, and Qwen have a strong correlation with human ratings, this greatly varies according to the items’ condition. In particular, the model scores for temporal anti-iconic and unrelated items have a much lower correlation with human ratings.
@@
We analyzed how the model size affects performance on PCD. We selected the Qwen model family, available in a wide range of sizes, and used the acceptability rating and perplexity tasks as a testbed. [^23] Few-shot and greedy search setup, i.e. the best for Qwen. Fig. 8 shows that Qwen’s performance and APS improve with model size, except for a slight drop for the latter in the largest variant. Nonetheless, the improvement rates are quite different. Performances seem to linearly scale with size, whereas the APS growth curve is flatter and eventually plateaus. The initial APS is markedly higher (2x) than the respective accuracy; accuracy and APS are nearly equal at the 14B mark; the accuracy of the 32B keeps improving while the APS is stale. A similar trend also holds for each relation [^24] Results for each are shown in App. K., despite showing greater variability. For example, the 0.5B variant is quite proficient with anti-iconic relations but almost incapable of modelling iconic ones. The model scale seems to correlate with smaller performance differences across relations.
@@ 6 Conclusion and Future Work
So far, the presented results reveal that, despite their increasing size, the knowledge of causal relations is still suboptimal in LLMs, which also show a strong tendency to confound temporal relations with causal ones, compounded with a limited abstraction from the surface linguistic presentation order of the events. We can conclude that, although LLMs have undoubtedly progressed along the ‘ladder of causation’ Pearl (2009), they have still several steps to climb.
@@ 7 Limitations
This study has several limitations that should be acknowledged.
One limitation of our dataset design is the deliberate removal of certain linguistic cues – such as anaphoric references and causal verbs – from the sentence pairs. While this may reduce the naturalness of the items (i.e., sentence pairs linked by a connective), retaining such cues could have allowed models to exploit shallow, surface-level heuristics rather than engaging in genuine causal reasoning, thereby introducing bias into the evaluation. Notably, the same sentence pairs (without pronouns) were also presented to human annotators. This design choice enabled a more controlled comparison, focusing on how both models and humans infer causal and temporal relations based on a deeper understanding of the events described.
The prompts used across different models were not specifically optimized for each of them. However, this decision was necessary to maintain the feasibility of our experiments and ensure a fair comparative evaluation among all models and tasks. Computational constraints played a significant role in shaping our methodology, influencing aspects such as batch size, model capacity, and the overall scope of the analysis. While we examined causal reasoning through various tasks, the selection of prompts and connectives represents only a subset of all potential analytical strategies.
Additionally, the number of closed-source models included in the analysis was limited, and the dataset size was relatively small, further constraining the generalizability of the findings. Experiments on model scaling were confined to the Qwen model, where we observed that larger models tend to equalize performance on tasks and competence. However, this observation requires further validation to confirm that it is not influenced by factors such as the specific prompt or model used, which would bolster the robustness of our claims.
Among the causal sentence pairs in ExpliCa, $50$ are labeled as ‘socially challenging’, indicating content that touches on sensitive or potentially offensive topics such as religion, abortion, immigration, gender identity, drug abuse, and bribery. We acknowledge that some sentences may be offensive to certain groups, but these themes were added to evaluate whether bias-mitigation strategies in LLMs would impact PCD performance. Due to space constraints, we plan to explore such aspects more in depth in future works.
Finally, we acknowledge that we focused only on LLMs and we did not include the so-called Large Reasoning Models (LRM) like OpenAI o1 OpenAI (2024b) or the very recent DeepSeek-R1 DeepSeek-AI (2025), which may have an advantage in such a task. This choice was mainly due to financial constraints, but they will be explored in future works.
@@ Acknowledgments
This work has been supported by i.) the PNRR MUR project PE0000013-FAIR (Spoke 1), funded by the European Commission under the NextGeneration EU programme; ii.) the EU EIC project EMERGE (Grant No. 101070918); iii.) the Italian Ministry of University and Research (MUR) in the framework of the PON 2014-2021 “Research and Innovation” resources – Innovation Action - DM MUR 1062/2021 - Title of the Research: “Modelli semantici multimodali per l’industria 4.0 e le digital humanities.”; and iv.) the GRF grant from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. PolyU 15612222).
@@
Appendix A Related Work Overview
Table 3 summarizes the main features that distinguish ExpliCa from other datasets on causal reasoning. We excluded from the overview the datasets based on structured data, benchmarks composed of existing datasets evaluating different aspects of causation, and the datasets included in ExpliCa.
Table 3: Overview of the feature belonging to existing datasets and ExpliCa.
Dataset References Frame Temporal Relation type No ling. cues Crowdsourced Domain
COPA Roemmele et al. 2011 Statements Implicit Commonsense
Causal TimeBank Mirza et al. 2014 Text Explicit News
CaTeRS Mostafazadeh et al. 2016 Text Implicit/Explicit Commonsense
Event Story Line Caselli and Vossen 2017 Text Implicit/Explicit News
BeCause Corpus Dunietz et al. 2017 Text Explicit Mostly News
TCR Ning et al. 2018 Text Implicit/Explicit News
FinCausal Mariko et al. 2022 Scenario Implicit/Explicit Finance
MECI Lai et al. 2022 Text Implicit/Explicit Wikipedia
CLadder Jin et al. 2023 Scenario Implicit Common/Anti/Nonsense
CRAB Romanou et al. 2023 Scenario Implicit/Explicit News
CausalNet Ashwani et al. 2024 Scenario Implicit/Explicit Mix
CausalBench Wang 2024 Scenario Implicit Mix
CausalProbe-2024 Chi et al. 2024 Scenario Implicit News
ExpliCa Our Work Statements Explicit Commonsense
Appendix B Dataset License
@@
Appendix C Human Annotation
@@
Appendix D Frequency Analysis
@@
Rare (00-$4$) 214 241 239 164 303
Uncommon ($5$-$56.4$) 244 244 295 174 282
Common ($56.5$-$498$) 173 205 272 169 380
Frequent ($499$-$7M$) 189 186 234 149 443
@@
Appendix E Models’ Output Cleaning
@@
Appendix F Prompt Engineering
@@ 3. Words:
Task Prompt Mistral Falcon Qwen Llama Gemma Avg
Accept. Rat. 0 9.99 23.2 14.7 13.74 18.85 16.10
1 12.23 24.38 21.76 16.14 32.39 21.38
2 15.59 25.05 14.37 17.7 38.81 22.30
3 12.81 29.52 20.89 20.09 31.34 22.93
4 14.62 29.86 18.25 17.81 30.62 22.23
Multiple-Choice 0 3.34 3.66 3.58 4.34 4.39 3.86
1 4.06 4.6 5.87 6.6 7.57 5.74
2 3.42 4.15 5.89 7.26 7.86 5.72
3 3.99 6.09 11.42 10.68 8.51 8.14
4 4.41 4.75 7.46 7.27 8.03 6.38
Cloze Test 0 13.41 15.52 20.41 14.2 21.38 16.98
1 9.03 12.42 24.07 20.05 19.98 17.11
2 9.87 12.06 16.64 13.39 15.02 13.40
3 12.72 13.91 21.18 21.34 23.73 18.58
4 13.52 14.93 17.65 12.13 14.95 14.64
Appendix G Models’ Specifications

### v4

@@ 4.1 Models
We selected $7$ generative LLMs. Specifically, we selected two open-weights models (Mistral-7B-Instruct-v0.3 and falcon-7b-instruct), three partially open models (Meta-Llama-3.1-8B-Instruct, gemma-2-9b-it, and Qwen2.5-7B-Instruct), and two commercial models, gpt4o and gpt4o-mini. Perplexity evaluation was not performed on commercial models as it is not permitted through the API. We used Qwen2.5 instruct models of different sizes (from 0.50.5B to 3232B parameters) for analyzing the impact of model scale. [^13] More details on models in App. F.
@@ 5 Results and Discussion
We carried out an in-depth analysis of several interesting aspects emerging from the experiments, as outlined in the following.
Core results are summarized in Fig. 3, [^14] Detailed results are shown in Tab. 5, App. G. which shows the average accuracy across all three prompting tasks for a specific model in a specific setting (e.g., Outlines in zero-shot). Violet bars represent the APS (a single numerical value) for each open model. These results are complemented by Tab. 2, which compares performances across prompting tasks, aggregated by model type.
The main trends observed are: i.) performance variability is high among models of the same scale in prompting tasks, despite similar APSs across open models; ii.) GPT4o outperforms all models, but its mini variant does not follow this trend and is still far from human performance; iii.) accuracy standard deviation shows that most open models exhibit inconsistent results across tasks, particularly with greedy search (Sec. 5.1); iv.) prompting performance is systematically lower than competence-level APSs in open models (Sec. 5.2).
5.1 Modeling PCD with Prompting
@@
[…]
Tab. 2 shows that, on average, PCD performed best when responses were framed using Outlines in a few-shot setting ($0.40$ overall, $0.61$ for GPT models). However, the highest task-specific accuracy was achieved in the acceptability rating task under a zero-shot setting with greedy search ($0.49$).
Moreover, the Outlines framework proved to be beneficial for all models on the multiple-choice task. Conversely, for the cloze test and acceptability rating task, performances either remained roughly the same or decreased substantially for both open models and GPTs. Strikingly, the top average accuracy of $0.61$ in Tab. 2 still reveals a strong gap with human data.
Fig. 3 shows that each model performs best under different settings. Each of GPT4o, Gemma, and Qwen achieved the highest mean accuracy ($0.72$, $0.53$, and $0.51$, respectively) in a distinct setup, surpassing GPT4o-mini ($0.50$). Standard deviation analysis (sd) indicates that task selection significantly affects smaller models’ performance, especially with greedy search (sd $0.2$ as in Tab. 2). Notably, GPT4o-mini exhibits a much wider performance range across both few-shot and zero-shot settings.
5.2 Modelling PCD with Perplexity
The APS results of the open models are shown in Fig. 3. The scores obtained by relying solely on perplexity are significantly higher ($0.63$ overall average) compared to those achieved in the prompting tasks. [^15] Perplexity values are in Tab. 8, App. G. Although Falcon performs poorly in the latter, it is the model with the best APS ($0.66$). By leveraging perplexity, the lowest APS too, reached by Qwen with $0.59$, surpasses the other open models in prompting tasks and GPT4o-mini as well. Although this result does not match GPT4o performances, it seems to confirm that models’ competence about causal relations encoded in their probabilistic predictions is more accurate than their prompting performances Hu and Levy (2023).
@@ 5.2 Modelling PCD with Perplexity
5.3 In-depth Analysis on ExpliCa
@@
First, we focus on the relation type between the events expressed in each dataset item, and their order. We report results on the acceptability rating prompting task, with zero-shot and greedy search (i.e., where the models performed best overall). [^16] All the models’ results, in all settings, are in App. G. We specifically focus on GPT4o and two open models, namely Gemma and Falcon, with the latter achieving the best APS. In Fig. 4, we see that models perform best on the causal relations in iconic order, with the exception of Falcon. [^17] Values are shown in Tab. 9, App. G. Worse performances are obtained for causal anti-iconic and temporal iconic, with the latter often being mistaken for causal iconic.
@@
An interesting pattern emerges by comparing APSs and Accuracy on acceptability ratings. APSs for a model are always higher than the respective acceptability ratings’ accuracy, except for temporal anti-iconic items. [^18] As for most of the models (see Tab. 8, App. G). However, by observing the models’ behavior, [^19] See App. H, Fig. 12 for the confusion matrices. we saw that both Gemma and Falcon tend to favor the temporal anti-iconic connective despite showing lower perplexity for the connective expressing a causal iconic relation (i.e., so). The performances of GPT4o do not seem to suffer the same biases toward specific cases. Nevertheless, intra-cases differences are up to 26%.
@@
We compared human ratings with model-generated ones in terms of their distribution and correlation. We show a comparison of models acceptability ratings with humans’ for the causal-temporal and unrelated subsets in Fig. 5.202020Answers without a rating are not shown in the distribution. We consider ratings of GPT4o and Gemma, [^21] Results for the zero-shot prompt with greedy search. These are the models with the highest accuracy for this task. and the (normalized) perplexity scores for Falcon, as it obtained the best APSs. A noticeable difference exists between human ratings of causal and temporal items vs. unrelated and model scores.
Fig. 6 shows Spearman correlation computed among human ratings and models’ acceptability ratings collected in zero-shot scenario. [^22] Individual results in Tab. 13, App. I. We observed that even though GPT4o, Gemma, and Qwen have a strong correlation with human ratings, this greatly varies according to the items’ condition. In particular, the model scores for temporal anti-iconic and unrelated items have a much lower correlation with human ratings.
@@
We analyzed how the model size affects performance on PCD. We selected the Qwen model family, available in a wide range of sizes, and used the acceptability rating and perplexity tasks as a testbed. [^23] Few-shot and greedy search setup, i.e. the best for Qwen. Fig. 7 shows that Qwen’s performance and APS improve with model size, except for a slight drop for the latter in the largest variant. Nonetheless, the improvement rates are quite different. Performances seem to linearly scale with size, whereas the APS growth curve is flatter and eventually plateaus. The initial APS is markedly higher (2x) than the respective accuracy; accuracy and APS are near equal at the 14B mark; the accuracy of the 32B keeps improving while the APS is stale. A similar trend holds true also for each relation, [^24] Results for each are shown in App. J. despite showing a higher variability. For example, the 0.5B variant is quite proficient with anti-iconic relations but almost incapable of modelling iconic ones. Model’s scale seem to correlate with less differences in performances among relations.
@@ 6 Conclusion and Future Work
So far, the presented results reveal that, despite their increasing size, the knowledge of causal relations is still suboptimal in LLMs, which also show a strong tendency to confound temporal relations with causal ones, compounded with a limited abstraction from the surface linguistic presentation order of the events. e can conclude that, although LLMs have undoubtedly progressed along the ‘ladder of causation’ Pearl (2009), they have still several steps to climb.
@@ 7 Limitations
This study has several limitations that should be acknowledged. The prompts used across different models were not specifically optimized for each of them. However, this decision was necessary to maintain the feasibility of our experiments and ensure a fair comparative evaluation among all models and tasks. Computational constraints played a significant role in shaping our methodology, influencing aspects such as batch size, model capacity, and the overall scope of the analysis. While we examined causal reasoning through various tasks, the selection of prompts, possible choices, and connectives represents only a subset of all potential analytical strategies. Additionally, the number of closed-source models included in the analysis was limited, and the dataset size was relatively small, further constraining the generalizability of the findings. Experiments on model scaling were confined to the Qwen model, where we observed that larger models tend to equalize performance on tasks and competence. However, this observation requires further validation to confirm that it is not influenced by factors such as the specific prompt or model used, which would bolster the robustness of our claims. Among the causal sentence pairs in ExpliCa, $50$ are labeled as ‘socially challenging’, indicating content that touches on sensitive or potentially offensive topics such as religion, abortion, immigration, gender identity, drug abuse, and bribery. We acknowledge that some sentences may be offensive to certain groups, but these themes were added to evaluate whether bias-mitigation strategies in LLMs would impact PCD performance. Due to space constraints, we plan to explore such aspects more in depth in future works. Finally, we acknowledge that we focused only on LLMs and we did not include the so-called Large Reasoning Models (LRM) like OpenAI o1 OpenAI (2024b) or the very recent DeepSeek-R1 DeepSeek-AI (2025), which may have an advantage in such a task. This choice was mainly due to financial constraints, but they will be explored in future works.
@@ Acknowledgments
This work has been supported by the PNRR MUR project PE0000013-FAIR (Spoke 1), by the EU EIC project EMERGE (Grant No. 101070918), and by the Italian Ministry of University and Research (MUR) in the framework of the PON 2014-2021 “Research and Innovation” resources – Innovation Action - DM MUR 1062/2021 - Title of the Research: “Modelli semantici multimodali per l’industria 4.0 e le digital humanities.”
@@
Appendix A Dataset License
@@
Appendix B Human Annotation
@@
Appendix C Frequency Analysis
@@
Rare (0-44) 214 241 239 164 303
Uncommon (55-$56.4$) 244 244 295 174 282
Common (56.556.5-$498$) 173 205 272 169 380
Frequent (499499-$7M$) 189 186 234 149 443
@@
Appendix D Models Answer Cleaning
@@
Appendix E Prompt Engineering
Task Prompt Mistral Falcon Qwen Llama Gemma Avg
Accept. Rat. 0 9.99 23.2 14.7 13.74 18.85 16.10
1 12.23 24.38 21.76 16.14 32.39 21.38
2 15.59 25.05 14.37 17.7 38.81 22.30
3 12.81 29.52 20.89 20.09 31.34 22.93
4 14.62 29.86 18.25 17.81 30.62 22.23
Multiple-Choice 0 3.34 3.66 3.58 4.34 4.39 3.86
1 4.06 4.6 5.87 6.6 7.57 5.74
2 3.42 4.15 5.89 7.26 7.86 5.72
3 3.99 6.09 11.42 10.68 8.51 8.14
4 4.41 4.75 7.46 7.27 8.03 6.38
Cloze Test 0 13.41 15.52 20.41 14.2 21.38 16.98
1 9.03 12.42 24.07 20.05 19.98 17.11
2 9.87 12.06 16.64 13.39 15.02 13.40
3 12.72 13.91 21.18 21.34 23.73 18.58
4 13.52 14.93 17.65 12.13 14.95 14.64
@@ 3. Words:
Appendix F Models Details